Nano Specialist is considering an upgrade project. The estimated cash flows from the upgrade project appear below. What is the project's payback period? Note that year 0 and year 1 cash flows are negative. (Answer in years, round to 2 places)

Year 0 cash flow = -85,000
Year 1 cash flow = -55,000
Year 2 cash flow = 11,000
Year 3 cash flow = 41,000
Year 4 cash flow = 30,000
Year 5 cash flow = 25,000
Year 6 cash flow = 23,000
Year 7 cash flow = 40,000
Answer #1

Cumulative cash flow for year 0 = -85,000

Cumulative cash flow for year 1 = -85,000 - 55,000 = -140,000

Cumulative cash flow for year 2 = -140,000 + 11,000 = -129,000

Cumulative cash flow for year 3 = -129,000 + 41,000 = -88,000

Cumulative cash flow for year 4 = -88,000 + 30,000 = -58,000

Cumulative cash flow for year 5 = -58,000 + 25,000 = -33,000

Cumulative cash flow for year 6 = -33,000 + 23,000 = -10,000

Cumulative cash flow for year 7 = -10,000 + 40,000 = 30,000

10,000 / 40,000 = 0.25

Payback period = 6 + 0.25 = 6.25 years